# Nucleon energy correlators as a probe of light-quark dipole operators at the EIC

### v1

@@ Appendix B Additional constraints on the dipole operators at the EIC
Figure 7: The projected double-operator constraints at HERA with $68\%$ C.L. on the real parts of the couplings, distinguishing between the up-channel (left column) and down-channel (right column). The top row shows constraints in the $\text{Re}[C_{q\gamma}]$-$\text{Re}[C_{qZ}]$ planes, while the bottom row shows the $\text{Re}[C_{qB}]$-$\text{Re}[C_{qW}]$ planes. All constraints are derived from the azimuthal asymmetries $A^{\sin\phi}_{UU}$ (purple) and $A^{\sin\phi}_{LU}$ (blue), assuming $\Lambda=1$ TeV. The red ellipses represent constraints from combining $A^{\sin\phi}_{UU}$ and $A^{\sin\phi}_{LU}$. The dashed purple and blue lines correspond to the single-operator constraints from $A^{\sin\phi}_{UU}$ and $A^{\sin\phi}_{LU}$ respectively.
The projected constraints on (${\rm Re}[C_{dB}]$, ${\rm Re}[C_{dW}]$) are presented in Fig. 5, and those on (${\rm Re}[C_{qZ}]$, ${\rm Re}[C_{q\gamma}]$) with $q=u,d$ are shown in Fig. 6. The latter couplings have been rescaled to dimensionless form via $C_{q\gamma(Z)}\equiv c_{q\gamma(Z)}/(v/\sqrt{2}\textrm{TeV}^{2})$. The most stringent constraints from combining $A_{UU}^{\sin\phi}$ and $A_{LU}^{\sin\phi}$ are highlighted in red, reaching values in the range $\mathcal{O}(0.01)\sim\mathcal{O}(0.1)$ for $C_{q\gamma}$. Individual limits from single-operator (double-operator) analyzes using $A_{UU}^{\sin\phi}$ and $A_{LU}^{\sin\phi}$ are indicated by purple and blue dashed lines (shaded regions), respectively. In some cases, the single-operator sensitivity from either $A_{UU}^{\sin\phi}$ or $A_{LU}^{\sin\phi}$ is too weak to be displayed; for instance, no limit is shown for ${\rm Re}[C_{dW}]$ from $A_{UU}^{\sin\phi}$. The similarity between the (${\rm Re}[C_{dZ}]$, ${\rm Re}[C_{d\gamma}]$) and (${\rm Re}[C_{uZ}]$, ${\rm Re}[C_{u\gamma}]$) constraints arises from their identical analytic dependence on the asymmetry expressions, as shown in Sec. A. Due to the relatively smaller transversity NEC for $d$-quark, the combined constraints on operators involving $d$-quark are approximately a factor of two weaker than those for $u$-quark operators.
@@
EIC AbdulKhalek:2021gbh 105 100 $70\%$ $[10,60]$ $[0.1,0.9]$ $[0.01,0.5]$
HERA Klein:2008di ; H1:2015ubc 318 0.4 $40\%$ $[30,150]$ $[0.1,0.9]$ $[0.01,0.5]$
LHeC LHeC:2020van 1300 50 $80\%$ $[100,1000]$ $[0.1,0.9]$ $[0.005,0.9]$
@@
Figure 8: The projected double-operator constraints at the LHeC with $68\%$ C.L. on the real parts of the couplings, distinguishing between the up-channel (left column) and down-channel (right column). The top row shows constraints in the $\text{Re}[C_{q\gamma}]$-$\text{Re}[C_{qZ}]$ planes, while the bottom row shows the $\text{Re}[C_{qB}]$-$\text{Re}[C_{qW}]$ planes. All constraints are derived from the azimuthal asymmetries $A^{\sin\phi}_{UU}$ (purple) and $A^{\sin\phi}_{LU}$ (blue), assuming $\Lambda=1$ TeV. The red ellipses represent constraints from combining $A^{\sin\phi}_{UU}$ and $A^{\sin\phi}_{LU}$. The dashed purple and blue lines correspond to the single-operator constraints from $A^{\sin\phi}_{UU}$ and $A^{\sin\phi}_{LU}$ respectively.
Figure 9: Comparison of projected (EIC, LHeC) and existing (HERA) constraints at $68\%$ C.L. on the real parts of the couplings, shown for the EIC (red), LHeC (blue, scaled by a factor of 1010), and HERA (green). The layout distinguishes between the up-channel (left column) and down-channel (right column). The top row corresponds to the $\text{Re}[C_{q\gamma}]$-$\text{Re}[C_{qZ}]$ planes and the bottom row corresponds to the $\text{Re}[C_{qB}]$-$\text{Re}[C_{qW}]$ planes. Each contour represents a combined constraint derived from the azimuthal asymmetries $A^{\sin\phi}_{UU}$ and $A^{\sin\phi}_{LU}$, assuming $\Lambda=1$ TeV.
Appendix C Comparison of constraints among HERA, the EIC and the LHeC
In addition to the EIC, this study analyzes the sensitivity of two other electron-proton colliders: HERA Klein:2008di ; Diaconu:2010zz and the Large Hadron-Electron Collider (LHeC) LHeC:2020van . HERA is included since our setup does not require nucleon polarization, while the LHeC represents a future high-energy frontier with enhanced sensitivity to new physics. For our numerical analyzes, we specifically consider the Run 6 phase of the LHeC. The parameters for these colliders are summarized in Table 1.

### v2

@@ Appendix E Additional constraints on the dipole operators at the EIC
EIC Abdul Khalek and others (2022) 105 100 $70\%$ $[10,60]$ $[0.1,0.9]$ $[0.01,0.5]$
HERA Klein and Yoshida (2008); Abramowicz and others (2015) 318 0.4 $40\%$ $[30,150]$ $[0.1,0.9]$ $[0.01,0.5]$
LHeC Agostini and others (2021) 1300 50 $80\%$ $[100,1000]$ $[0.1,0.9]$ $[0.005,0.9]$
@@
Figure 8: The projected double-operator constraints at HERA with $68\%$ C.L. on the real parts of the couplings, distinguishing between the up-channel (left column) and down-channel (right column). The top row shows constraints in the $\text{Re}[C_{q\gamma}]$-$\text{Re}[C_{qZ}]$ planes ($C_{q\gamma(Z)}\equiv c_{q\gamma(Z)}/(v/\sqrt{2}\textrm{TeV}^{2})$), while the bottom row shows the $\text{Re}[C_{qB}]$-$\text{Re}[C_{qW}]$ planes. All constraints are derived from the azimuthal asymmetries $A^{\sin\phi}_{UU}$ (purple) and $A^{\sin\phi}_{LU}$ (blue), assuming $\Lambda=1$ TeV. The red ellipses represent constraints from combining $A^{\sin\phi}_{UU}$ and $A^{\sin\phi}_{LU}$. The dashed purple and blue lines correspond to the single-operator constraints from $A^{\sin\phi}_{UU}$ and $A^{\sin\phi}_{LU}$ respectively.
The projected constraints on (${\rm Re}[C_{dB}]$, ${\rm Re}[C_{dW}]$) are presented in Fig. 6, and those on (${\rm Re}[C_{qZ}]$, ${\rm Re}[C_{q\gamma}]$) with $q=u,d$ are shown in Fig. 7. The latter couplings have been rescaled to dimensionless form via
$C_{q\gamma(Z)}\equiv c_{q\gamma(Z)}/(v/\sqrt{2}\textrm{TeV}^{2})~.$ (38)
The most stringent constraints from combining $A_{UU}^{\sin\phi}$ and $A_{LU}^{\sin\phi}$ are highlighted in red, reaching values in the range $\mathcal{O}(0.01)\sim\mathcal{O}(0.1)$ for $C_{q\gamma}$. Individual limits from single-operator (double-operator) analyses using $A_{UU}^{\sin\phi}$ and $A_{LU}^{\sin\phi}$ are indicated by purple and blue dashed lines (shaded regions), respectively. In some cases, the single-operator sensitivity from either $A_{UU}^{\sin\phi}$ or $A_{LU}^{\sin\phi}$ is too weak to be displayed; for instance, no limit is shown for ${\rm Re}[C_{dW}]$ from $A_{UU}^{\sin\phi}$. The similarity between the (${\rm Re}[C_{dZ}]$, ${\rm Re}[C_{d\gamma}]$) and (${\rm Re}[C_{uZ}]$, ${\rm Re}[C_{u\gamma}]$) constraints arises from their identical analytic dependence on the asymmetry expressions, as shown in Sec. D. Due to the relatively smaller transversity NEC for $d$-quark, the combined constraints on operators involving $d$-quark are approximately a factor of two weaker than those for $u$-quark operators.
Figure 9: The projected double-operator constraints at the LHeC with $68\%$ C.L. on the real parts of the couplings, distinguishing between the up-channel (left column) and down-channel (right column). The top row shows constraints in the $\text{Re}[C_{q\gamma}]$-$\text{Re}[C_{qZ}]$ planes ($C_{q\gamma(Z)}\equiv c_{q\gamma(Z)}/(v/\sqrt{2}\textrm{TeV}^{2})$), while the bottom row shows the $\text{Re}[C_{qB}]$-$\text{Re}[C_{qW}]$ planes. All constraints are derived from the azimuthal asymmetries $A^{\sin\phi}_{UU}$ (purple) and $A^{\sin\phi}_{LU}$ (blue), assuming $\Lambda=1$ TeV. The red ellipses represent constraints from combining $A^{\sin\phi}_{UU}$ and $A^{\sin\phi}_{LU}$. The dashed purple and blue lines correspond to the single-operator constraints from $A^{\sin\phi}_{UU}$ and $A^{\sin\phi}_{LU}$ respectively.
Appendix F Comparison of constraints among HERA, the EIC and the LHeC
In addition to the EIC, this study analyzes the sensitivity of two other electron-proton colliders: HERA Klein and Yoshida (2008); Diaconu et al. (2010) and the Large Hadron-Electron Collider (LHeC) Agostini and others (2021). HERA is included since our setup does not require nucleon polarization, while the LHeC represents a future high-energy frontier with enhanced sensitivity to new physics. For our numerical analyses, we specifically consider the Run 6 phase of the LHeC. The parameters for these colliders are summarized in Table 1.
@@ Appendix F Comparison of constraints among HERA, the EIC and the LHeC
Figure 10: Comparison of projected (EIC, LHeC) and existing (HERA) constraints at $68\%$ C.L. on the real parts of the couplings, shown for the EIC (red), LHeC (blue, scaled by a factor of 1010), and HERA (green). The layout distinguishes between the up-channel (left column) and down-channel (right column). The top row corresponds to the $\text{Re}[C_{q\gamma}]$-$\text{Re}[C_{qZ}]$ planes ($C_{q\gamma(Z)}\equiv c_{q\gamma(Z)}/(v/\sqrt{2}\textrm{TeV}^{2})$) and the bottom row corresponds to the $\text{Re}[C_{qB}]$-$\text{Re}[C_{qW}]$ planes. Each contour represents a combined constraint derived from the azimuthal asymmetries $A^{\sin\phi}_{UU}$ and $A^{\sin\phi}_{LU}$, assuming $\Lambda=1$ TeV.
Appendix G Discussions on the uncertainties from non-perturbative inputs
In this section, we evaluate the impact of theoretical uncertainties associated with the nonperturbative inputs on our projected constraints. The transversity NEC, $h_{1}^{t}$, which serves as the crucial input for our asymmetry calculation, is estimated with the first transverse moment of the Boer-Mulders quark TMD function, $\mathscr{h}_{1}^{\perp q}(x,\bm{k}_{\perp}^{2})$. Consequently, uncertainties in the extraction of the Boer-Mulders TMD propagate to our final constraints on the dipole couplings. For comparison and simplicity, we adopt the scale evolution implemented in the TMD fits for the corresponding NEC moments in this section.
To assess the robustness of our results, we consider two factors affecting the nonperturbative uncertainties: (1) the uncertainties within the Barone et al. fit Barone et al. (2010) we adopted in our analysis, and (2) the impact of using different Boer-Mulders TMD fits.
G.1 Uncertainties from the Barone et al. Fit
In our primary analysis, we utilized the parameterization of the Boer-Mulders TMD provided by Barone et al. Barone et al. (2010). This fit is based on a combined analysis of COMPASS and HERMES data and relies on the ansatz that the Boer-Mulders function is proportional to the Sivers function, $f_{1T}^{\perp q}$:
$\mathscr{h}_{1}^{\perp q}(x,\bm{k}_{\perp}^{2})=\lambda_{q}\mathscr{f}_{1T}^{\perp q}(x,\bm{k}_{\perp}^{2}).$ (39)
The dominant uncertainty in this extraction resides in the flavor-dependent normalization factors, $\lambda_{q}$. The best-fit values and their associated uncertainties are given by:
$\lambda_{u}=2.0\pm 0.1,\quad\lambda_{d}=-1.111\pm 0.001.$ (40)
To estimate the impact of these uncertainties on our projected constraints, we generate an envelope with the upper and lower values of $\lambda_{q}$. We demonstrate the effects of this envelope on the projected constraints in Fig. 11 (Left). The variation in the TMD parameters leads to only a slight broadening of the allowed region, indicating that the uncertainties in this fit do not significantly affect the competitiveness of our projected bounds.
Figure 11: Left: Constraints based on the fit of Barone et al., including the associated uncertainties. Right: Comparison of the constraints on the dipole couplings obtained using different fits of the Boer–Mulders TMD as input (Here, we have used the central values).
G.2 Impact of different phenomenological fits
To assess the impact of different phenomenological fits, we compared the constraints derived using the Barone et al. fit with those obtained from two other independent parameterizations of Boer-Mulders TMD available in the literature:
  • Christova et al. Christova et al. (2020): A more recent extraction based on COMPASS SIDIS data. Notably, this analysis does not assume the proportionality ansatz between the Boer–Mulders and Sivers functions used in Barone et al..
  • Zhang et al. Zhang et al. (2008): An extraction utilizing data from unpolarized $p+D$ Drell-Yan processes, providing a complementary constraint from a different scattering channel.
For this comparison, we adopted the central best-fit values from each parameterization. The resulting constraints are displayed in Fig. 11 (Right).
We observe consistent bounds derived from all three parameterizations, indicating our results are not sensitive to the choice of nonperturbative inputs. This insensitivity can be understood as follows. The asymmetry is sensitive primarily to the first transverse moment of the Boer–Mulders function, weighted by $\bm{k}_{\perp}^{2}/M_{N}^{2}$. While the detailed functional forms of $\mathscr{h}_{1}^{\perp q}(x,\bm{k}_{\perp}^{2})$ differ among the fits, they predict similar effective transverse widths. Furthermore, our analysis involves an integration over a wide range of Bjorken-$x$, which tends to wash out local differences in the $x$-dependence of the various parametrizations.
The above analysis demonstrates that the projected sensitivity to the light-quark dipole couplings is stable against the current theoretical uncertainties associated with the nonperturbative Boer–Mulders TMD inputs.
Appendix H Comparison with other collider probes on light-quark dipole operators
In this section, we provide a detailed comparison of our projected sensitivities with existing or proposed constraints from other collider probes. We focus specifically on the constraints derived from existing Drell-Yan data at the LHC Boughezal et al. (2021) and from a proposed dihadron measurement at the EIC Wen et al. (2024a), as these represent the most relevant competing methods for probing light-quark dipole moments at colliders.
H.1 Comparison with Drell-Yan constraints at the LHC
We first compare our projections with constraints derived from the Drell-Yan process ($pp\to Z/\gamma^{*}(\to\ell^{+}\ell^{-})+X$) at the LHC Boughezal et al. (2021). As discussed in our manuscript, in the Drell-Yan process, the contribution from dimension-6 dipole operators appears only at $\mathcal{O}(\Lambda^{-4})$ in the cross-section, due to the chiral structure of the interaction. This leads to two limitations. First, the signal is suppressed relative to our observable, which arises from the $\mathcal{O}(\Lambda^{-2})$ interference between the dipole operator and the SM amplitude. This suppression, however, can be compensated by the higher energy of the LHC, compared to the EIC and HERA in our proposal. Second, the observables in Ref. Boughezal et al. (2021) suffer from contamination by various other SMEFT operator contributions. Specifically, non-dipole dimension-6 operators can contribute at $\mathcal{O}(\Lambda^{-2})$ and significantly dominate over the dipole terms, while dimension-8 operators also enter at $\mathcal{O}(\Lambda^{-4})$. Consequently, deriving robust constraints from Drell-Yan requires the assumption that these contaminating operators are absent. In contrast, our observable is uniquely sensitive to the dipole–SM interference at the leading $\mathcal{O}(\Lambda^{-2})$ order. This feature not only enhances sensitivity but also protects the analysis from contamination by other potential UV effects. Furthermore, because the Drell-Yan cross-section depends on the squared modulus of the couplings ($|C|^{2}$), it cannot resolve the complex phase of the dipole operators. Our method allows for the separate extraction of the real and imaginary parts via $\cos\phi$- and $\sin\phi$-weighted asymmetries.
[…]
In Fig. 12, we show the quantitative comparison. We include the reported bounds from Ref. Boughezal et al. (2021) (rescaled to $68\%$ C.L.) alongside our projections for both the conservative ($\mathcal{L}=100~\mathrm{fb}^{-1}$) and high-luminosity ($\mathcal{L}=1000~\mathrm{fb}^{-1}$) EIC configurations. Even with the conservative luminosity, our observable yields constraints stronger than the current LHC Drell-Yan bounds. With the high-luminosity configuration, the sensitivity is further improved, providing a powerful and theoretically clean probe of the dipole couplings.
H.2 Comparison with the dihadron approach for the EIC
Ref. Wen et al. (2024a) (Wen et al.) achieves a nonzero dipole-SM interference by observing collinear dihadron production in the DIS with an unpolarized proton beam ($ep\to e(h_{1}h_{2})X$).
A direct comparison between the results in Ref. Wen et al. (2024a) and our main manuscript is not straightforward due to differences in the assumed luminosity ($\mathcal{L}=1000\ \mathrm{fb}^{-1}$ in Ref. Wen et al. (2024a) vs. $\mathcal{L}=100\ \mathrm{fb}^{-1}$ in our baseline) and operator conventions. To facilitate a like-for-like comparison, we have performed a new analysis of our observable assuming the high-luminosity configuration ($\mathcal{L}=1000\ \mathrm{fb}^{-1}$) and converted the results of Ref. Wen et al. (2024a) into our operator convention using the relation:
$C_{q\gamma}=\frac{e}{\sqrt{2}}\left(\frac{1\,\mathrm{TeV}}{v}\right)^{2}\Gamma^{q}_{\gamma}~,$ (41)
where $v\approx 246$ GeV is the Higgs vacuum expectation value, and $\Gamma^{q}_{\gamma}$ is the dipole coupling in the convention of Ref. Wen et al. (2024a).
Figure 13: A comparison between the projected constraints reported in Ref. Wen et al. (2024a) based on dihadron production at the EIC and our corresponding results. Here, we have adopted the same integrated luminosity, ${\cal L}=1000~\mathrm{fb}^{-1}$, for a meaningful comparison.
In Fig. 13, we compare the combined constraints on $\mathrm{Re}[C_{u\gamma}]$ and $\mathrm{Re}[C_{d\gamma}]$. As shown in the figure, our observable yields a significantly smaller allowed region, excluding a substantial portion of the parameter space allowed by the dihadron analysis. While both our observables and Ref. [18] exhibit strong correlations between up- and down-quark dipole couplings, they yield different flat directions. Consequently, a combined analysis could substantially mitigate these correlations. Additionally, under a single-operator assumption, Ref. Wen et al. (2024a) estimates the magnitude of the constraint on $C_{q\gamma}$ (after conversion to our convention) to be approximately $0.1$. Our projected sensitivities for the same luminosity improve upon this estimation by approximately one order of magnitude.
Finally, we note that the observable proposed in this work offers distinct experimental advantages. It requires neither the polarized proton beams utilized in Ref. Boughezal et al. (2023) nor the semi-inclusive measurements involving particle identification and multi-particle tracking required for the dihadron method in Ref. Wen et al. (2024a).